OpenStudy (anonymous):

Convert each rate using analysis. Round to the nearest tenth if necessary.45 mi/h = ___ ft/s

OpenStudy (cwrw238):

first convert miles to feet by multiplying by 5280 now divide the result by the number of seconds in an hour

OpenStudy (anonymous):

1mile=1760 yards 1 yard=3feet 1 hour=60minute=60*60=3600 seconds \[ 1mile per hour =\frac{ 1*1760*3 }{ 1*3600 } feet / \sec\]